OpenStudy (anonymous):

Martha drives to work every day and passes two independently operated traffic lights. The probability that both lights are green is 0.31. The probability that the first light is green is 0.51. What is the probability that the second light is green, given that the first light is green? (5 points) 0.61 0.65 0.78 0.73

OpenStudy (kropot72):

Let the first light be A and the second B. P(A green) and P(B green) = P(A green) * P(B green) = 0.31 P(A green) = 0.51 Therefore P(B green) given A is green is: \[\frac{0.31}{0.51}\approx0.61\]

OpenStudy (anonymous):

Ty so much!!

OpenStudy (anonymous):

This was correct!!